### v1

@@ 1 Introduction
Figure 1: Research gaps and the key question in urban visual perception studies, followed by our methodology to bridge them.
@@ 2.1 Demographics and personalities as moderators of perception
Figure 2: Statistical difference of perception scores by demographics. Perception Q scores are calculated from ratings by participants in each subgroup for all (without location grouping) and each location. Locations with no significant differences in any indicator and demographic groups with fewer than four samples are not shown. Significance thresholds *$p<0.05$.
@@
We evaluated whether the mean SVI perception Q scores significantly differ across participants grouped by demographic characteristics and personality traits. We found differences between demographic and personalities subgroups but they differ by locations, i.e., where participants are from, and perceptual indicators (Figure 2). The highest number of significant differences happened in populations grouped by their age group, and across all five demographics (gender, age group, annual household income bracket, education level, and race and ethnicity) our new indicator walk had the most intra-country differences. The lowest number of significant differences were found in negative indicators, none for boring and one for depressing, across all demographics in all countries. In terms of countries, we found differences in participants from San Francisco and Santa Clara (USA) in four out of the five demographic characteristics and no differences in any demographic group of participants from the greater Amsterdam region (Netherlands). We noted gender differences under the live nearby and cycle indicators for specific countries (Nigeria and USA, respectively), and across all locations (e.g., Q scores calculated without grouping by countries) under the safe indicator (top plot in Figure 2).
@@
Differences between age groups occurred mostly among participants in the “21-29” and “40-49” groups, “40-49” and “Above 50” groups, and “30-39” and “40-49” groups (second plot from the top in Figure 2). The first pair of groups has the largest gap between them and leads to differences in perceptions for the beautiful and green indicators in Singapore and walk in Nigeria. The second age group encompasses all participants above 40 years old. This led to perception differences for the depressing indicator in Chile and walk in Nigeria. Lastly, the middle pair of groups of participants, between 30 and 49 years old, had perception differences for the walk indicator in Chile. In groups formed by Annual Household Income (AHI) brackets, we noticed the difference between participants in the “Bottom” bracket in relation to the “Middle” and “Upper” brackets for perception scores in the safe indicator in the USA. Differences were also found in Chile between participants in the “Bottom” and “Middle” bracket in the wealthy indicator (third plot from the top in Figure 2).
Differences among participants’ education levels were only found in participants from the USA, primarily the ones with a bachelor’s degree. These differences happened between bachelor’s and postgraduate degree holders and between bachelor’s degree holders and participants with a college, vocational training, or diploma, for the safe and lively indicators, respectively (fourth plot from the top in  Figure 2).
@@
Lastly, we found differences in groups determined by the participants’ self-reported race and ethnicity in the USA and Chile (bottom plot in Figure 2). The former shows differences between “Asian or Pacific Islanders” and “White or Caucasian” participants for the lively indicator, “Asian or Pacific Islanders” and “Hispanic and Latino” as well as “Hispanic and Latino” and “White or Caucasian” for the cycle indicator. In Chile, we also found differences between “Hispanic and Latino” and “White or Caucasian” for the walk indicator. The demographic groups on our dataset are consistent with the population mix in the sampled cities of San Francisco and Santa Clara (USA) and Santiago (Chile).
@@
Figure 3: Statistical difference of perception scores by personality traits. Perception Q scores are calculated from ratings by participants in each subgroup for all (without location grouping) and each location. Locations with no significant differences in any indicator are not shown. Significance threshold *$p<0.05$.
@@
Perceptual differences among participants with high and low personality scores in extraversion exist for the lively and beautiful indicators in Nigeria and San Francisco and Santa Clara (USA). Among participants with agreeableness traits, we found differences in the new green indicator in Singapore and San Francisco and Santa Clara (USA). As mentioned earlier, we found the most differences between participants with conscientiousness personality traits. These differences are notable for the safe and beautiful indicators in participants from San Francisco and Santa Clara (USA) and the greater Amsterdam region (Netherlands), respectively. The differences in the new indicator walk, however, exist in participants from Santiago (Chile) and among all combined participants. Lastly, we saw only one difference in each personality trait of neuroticism and openness, for the wealthy and boring indicators in Singapore and Nigeria, respectively.
@@ 2.5 Perceptual indicators correlation
We performed a correlation analysis and found that our new proposed indicators correctly capture the positive and negative sentiment of the existing perception indicators. The new indicators live nearby, walk, and cycle show a positive correlation (R$>$0.40) with the traditional positive indicators (safe, lively, wealthy, and beautiful) and a negative correlation (R$<$-0.40) with the negative indicators (boring and depressing) (Figure 5). An exception is seen in the relationship between cycle and lively (R$>$0.20) and cycle and boring (R$<$-0.20). The newly proposed green indicator shows an inverse result of negative correlation with a positive indicator (lively) and vice versa (boring).
We further investigated these relationships with polynomial regressions. Figure 10 shows the best fitting curve (linear, quadratic, or cubic), in terms of $R^{2}$, for the different perceptual pairs. The green perceptual indicator seems to best correlate linearly with safe and lively, R values of 0.30 and -0.19, respectively. There is a weak relationship between the green perceptual indicator and wealthy, boring, and depressing indicators ($0.03\leq R^{2}\leq 0.05$). Nevertheless, the relationship between green and beautiful, although relatively low ($R^{2}=0.12$), is as much as three times higher. An extended correlation analysis is shown in Appendix EFigure 9.
@@ 3 Discussion
Across all cities chosen, the statistical analysis revealed that differences in global urban visual perception are explained by demographics and personalities (Figure 2). While existing work found safety perceptions, in specific locations, are explained by gender [34, 35, 15, 14, 16], our findings revealed a potential aggregated difference when considering all five locations together. Similar to work done in [11], we did not find statistically significant differences between age groups for safety perceptions, although the age group representation in said work was less granular: below or above 50 years old. Another work also did not find significant differences in perception scores by age or gender [9], although similarly to [11], age groups were analyzed as below or above a given age, i.e., 28 years old in said study. Unlike work in [34], we did not find significant differences in the boring indicator for gender, or any, demographic subgroups. Nevertheless, our findings show that in other indicators, both traditional and new, there are significant differences for all five demographic characteristics in different countries. More importantly, these significant differences are found in cities from continents often overlooked by the research community, e.g., South America and Africa [36].
@@
Our findings on personality traits follow this same trend (Figure 3). While there are fewer significant differences compared to those observed across demographic subgroups, Santiago (Chile) and Abuja (Nigeria) exhibit differences among residents with distinct personality traits, e.g., personality scores in either Q1 or Q3 of their cities. Personality traits have been linked with subjective well-being, for example, conscientiousness was found to be a good predictor of life satisfaction [37]. Personality subgroups based on this trait led to the most significant differences among most countries, even when all participants’ ratings were combined (middle plot in Figure 3). These results open the opportunity to quantify how urban streetscapes are perceived by residents of different levels of well-being. The remaining four personality traits accounted to fewer significant associations, at most half of the ones found within conscientiousness groups. Nevertheless, our results reinforce ample literature correlating personality and landscape preference and the relationship of personality and aesthetic judgment [38].
@@
Our quantitative study expands the understanding of machine learning model magnitude bias on urban visual perception prediction, corroborates the perception differences from different populations and how different cities are perceived, and quantifies the influence of someone’s location when rating their cities (Figure 4). Supporting previous ideas of cities’ identity and their characteristics being captured by machine learning models [39, 40], we found that the predictions of ViT-PP2, a model trained on PP2 [6], differ consistently with our diverse dataset. While we noted differences between how participants from each city rated all SVI and between how all participants rated each city (top and middle plot in Figure 4), the predicted perception scores paint optimistic and pessimistic scenarios for the different cities. In both scenarios, albeit with some exceptions, the predicted perceptions of a one-size-fits-all and off-the-shelf model are statistically higher than the ground truth scores under positive indicators and statistically lower under negative indicators. Moreover, participants seem to rely on visual familiarity, influenced by cultural differences and location-specific norms, when rating their own city and others. Despite not having information on the location of the SVI, i.e., the pairwise comparison did not show the city the SVI is from, participants appeared to either be familiar with their city or could identify characteristics similar to their city and use that as a reference when rating other cities they might not be familiar with (bottom plot in Figure 4).
@@
Our work asserts the importance of multi-city and multi-population analysis. Previous multi-country work either analyzed only a single country of residency of a dataset [9], did not consider participants’ location [6], or focused on a single feature, e.g., biophilia [8]. Equally importantly, our work evidences that targeted intervention should consider the perception of locals in their own city [42], as it will differ greatly if the perceptual evaluation is done using external SVI or by a machine learning model prediction [16]. Nevertheless, we do not claim generalization or universality of these patterns, and we are limited to evaluating a single city in each country as the scope of this work. While we handled inter-country biases by analyzing demeaned perception scores in their respective locations (bottom plot in Figure 4), in-depth analysis on cross-cultural, location-based biases remains a promising research direction. At the same time, intra-country differences should be further analyzed, especially in larger and more diverse countries.
@@
Despite the demographic diversity and geographic location of participants, most of our proposed indicators (live nearby, cycle, and walk) capture the trends of the existing six indicators. These new dimensions are able to expand perception studies with indicators reflecting a user’s willingness to be associated with an urban scene (e.g., live, walk, or cycle in the surroundings). The green indicator introduces the opportunity to expand traditional pixel ratio greenery metrics, e.g., Green View Index (GVI), with subjective human perception data. While the relationships found with other perceptual indicators were more complex and nonlinear, we found around half of the statistically significant differences in perception scores between demographic characteristics (9 out of 18) and personality traits (4 out of 10) under these four new indicators. These highlight the role of these indicators, and we propose that researchers consider including them in their studies where possible and relevant.
@@
This work contributes to the few studies in multi-city perception, emphasizes the importance of human-centric data collection, and releases a global and diverse perception dataset. Methodologically, we introduced a global and, participants-wise, demographically balanced urban visual perception called SPECS and we thoroughly analyzed perception scores by demographic and personality subgroups. Moreover, we quantified and statistically verified, for the first time, the magnitude bias of off-the-shelf machine learning models in perceptual predictions despite being trained on a global dataset like PP2, and introduced and validated new indicators relevant to the six commonly used. Our results draws us closer to understanding in what contexts, e.g., locations and perceptual indicators, urban visual perception is most affected by demographics and personality. While we are not able to pinpoint a specific number or percentage of how much people’s profiles drive perception responses, we found that as such it can lead to significant differences depending on the profiles and locations. These insights on profiles and location subgroups can be extended to other urban studies, such as building exterior evaluation for urban development and strategies [43]. Street-level image-based indicators closely correlate with a city’s social structure and economic development [44]. Thus, as SVI captures the human-level perspective, more closely aligned with people’s experiences in cities, studying and understanding urban streetscapes allows researchers and stakeholders to take informed actions [45]. While PP2 remains an impressive effort valuable for urban science, our analysis reveals its limitations in localized contexts, suggesting that generalized or one-size-fits-all models may produce distorted results. We recommend that researchers fine-tune their perception prediction models with local data, even though this may require additional data collection efforts. This approach is likely to produce more accurate and better performing models [11, 16].
@@
Secondly, we acknowledge the limited number of cities and countries chosen. Our approach of one city per country per continent limits the generalizability of our findings to the broader urban landscape worldwide. Rather than claiming global or country-level representativeness, our study focuses on determining whether perceptual indicators vary systematically across demographic groups in diverse urban environments. To mitigate these limitations, we employed stratified sampling of both urban environments (image selection) and participant demographics, ensuring proportional representation of local socioeconomic contexts within each selected city.
@@
Future research should prioritize analyzing interactions between multiple demographic factors, e.g., older adults with different walk abilities [46], and their combined effects on various perceptual indicators. As our study nested participant subgroups by demographics, the resulting smaller participant pools yielded fewer ratings per perceptual indicator (e.g., safety perception scores based solely on older women’s ratings). To address this limitation, subsequent studies could either increase participant numbers or ratings per image. The latter could be achieved by reducing the number of images per city, though potentially compromising urban landscape diversity, or by focusing on fewer perceptual indicators, such as exclusively studying safety perception.

### v2

@@ 1 Introduction
Figure 1: Research gaps and key questions in urban visual perception studies, followed by our methodology to bridge them. The profile survey included demographics (gender, age group, education level, annual household income, and race and ethnicity) and personality traits (extraversion, agreeableness, conscientiousness, neuroticism, openness). The human perception ratings included the traditional six indicators (safe, wealthy, lively, beautiful, depressing, and boring) and new proposed four indicators (live nearby, walk, cycle, and green) for pairwise comparisons.
@@ 2.1 Demographics and personalities as moderators of perception
Figure 2: Statistical difference of perception scores by demographics. Perception Q scores are calculated from ratings by participants in each subgroup for all (without location grouping) and each location. Locations with no significant differences in any indicator and demographic groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per subgroup) are not shown. Significance thresholds *$p<0.05$.
@@
We evaluated whether the mean SVI perception Q scores significantly differ across participants grouped by demographic characteristics and personality traits. We found differences between demographic and personalities subgroups but they differ by locations, i.e., where participants are from, and perceptual indicators (Figure 2). The highest number of significant differences happened in populations grouped by gender, and across all five demographics (gender, age group, annual household income bracket, education level, and race and ethnicity) the traditional indicator safe had the most intra-country differences, followed by our new indicator cycle. No significant differences were found in negative indicators and only one significant difference was found in the live nearby and lively indicator in gender and race and ethnicity subgroups, respectively, in the USA. In terms of countries, we found differences in participants from San Francisco and Santa Clara (USA) in four out of the five demographic characteristics and no differences in any demographic group of participants from Santiago (Chile) or Singapore. We noted gender differences under the live nearby, and cycle indicators for specific countries (USA and Nigeria, respectively), and across all locations (e.g., Q scores calculated without grouping by countries) under the safe indicator (top plot in Figure 2).
@@
Differences between age groups occurred mostly among participants in the “21-29” and “Above 50” groups (second plot from the top Figure 2). This group had differences in perception for the cycle indicator in Netherlands. In groups formed by Annual Household Income (AHI) brackets, we noticed the difference between participants in the “Bottom” bracket in relation to the “Middle” and “Upper” brackets for perception scores in the safe indicator in the USA (middle plot Figure 2).
@@
Differences among participants’ education levels were only found in participants from the USA. These differences happened between bachelor’s and postgraduate degree holders for the safe indicator (fourth plot from the top  Figure 2). Lastly, we found differences in groups determined by the participants’ self-reported race and ethnicity in the USA (bottom plot in Figure 2). The differences for the lively indicator appeared between “Asian or Pacific Islanders” and “White or Caucasian” participants. The demographic groups on our dataset are consistent with the population mix in the sampled cities of San Francisco and Santa Clara (USA)
While overall we found fewer differences within demographic interaction subgroups, i.e., a total of four compared to eight in single demographic groups (Figure 2), among all the combinations of ‘gender $\times$ age group $\times$ AHI’ subgroups, half the differences happened within-country and half across all locations (Figure 9). Differences formed by all three demographics combined were found for the lively indicator between “Female x 30-39 x Bottom” and “Male x 21-29 x Middle” subgroups only when all participants’ responses were combined, although under a small sample size ($n\geq 5$, top plot in Figure 9). Within ‘gender $\times$ age group’ subgroups perception differences for the walk indicator between genders of the “Above 50” age group in the USA also under a smaller sample size ($n\geq 5$, second plot from the top Figure 9). Then, differences for the safe indicator between genders of the “Middle” AHI group were found in Netherlands under a larger sample size. Finally, perceptual differences between age groups “21-29” and “30-39” of the “Bottom” AHI bracket were found for the lively indicator across all locations. Both perceptual differences for safe and lively were found with higher sample size ($n\geq 20$, last two bottom plots Figure 9, respectively).
@@
Figure 3: Statistical difference of perception scores by personality traits. Perception Q scores are calculated from ratings by participants in each subgroup for all (without location grouping) and each location. Locations with no significant differences in any indicator and personality groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per subgroup) are not shown. Significance threshold *$p<0.05$.
@@
Perceptual differences among participants with high and low personality scores in extraversion exist for the lively and beautiful indicators in Nigeria and San Francisco and Santa Clara (USA) (top plot Figure 3). Among participants with agreeableness traits, we found differences in the new green indicator in Singapore (seconc plot from the top Figure 3). Differences between participants with conscientiousness personality traits are notable for the safe and walk indicators in participants from San Francisco and Santa Clara (USA) and across all locations, respectively (middle plot Figure 3). Lastly, we saw one difference in the personality trait of neuroticism for the wealthy indicator in Singapore (fourth plot from the top Figure 3).
@@ 2.5 Perceptual indicators correlation
We performed a linear correlation analysis and found that our new proposed indicators correctly capture the positive and negative sentiment of the existing perception indicators. The new indicators live nearby, walk, and cycle show a positive correlation (R$>$0.40) with the traditional positive indicators (safe, lively, wealthy, and beautiful) and a negative correlation (R$<$-0.40) with the negative indicators (boring and depressing) (Figure 5). An exception is seen in the relationship between cycle and lively (R$>$0.20) and cycle and boring (R$<$-0.20). The newly proposed green indicator shows an inverse result of negative correlation with a positive indicator (lively) and vice versa (boring). An extended linear correlation analysis is shown in Appendix EFigure 10.
We further investigated non-linear relationships with polynomial regressions. Figure 12 shows the best fitting curve in red (linear, quadratic, or cubic) in terms of $R^{2}$, for the different perceptual pairs, i.e., new indicator vs traditional indicator. In most instances, the best fitting model for live nearby, walk, and cycle is a linear model (in three, five, and four out of the six perceptual pairs, respectively). In the remaining perceptual pairs, the best fitting model has, in all but one case, an $R^{2}$ value 0.1 higher than the respective linear model, with the highest difference being 0.2 in the live nearby vs wealthy pair. This suggests that these three new perceptual indicators are predominately linearly correlated with the traditional six indicators. The green perceptual indicator appears to have a more complex relationship with the traditional six indicators (12(b)). This indicator best correlates linearly with safe and lively, R values of 0.30 and -0.19, respectively and $R^{2}$ values of 0.09 and 0.04. There is a weak non-linear relationship between the green perceptual indicator and wealthy, boring, and depressing indicators ($0.03\leq R^{2}\leq 0.05$), though still three times higher than $R^{2}$ values from linear models. Nevertheless, the relationship between green and beautiful, although relatively low ($R^{2}=0.12$ with a cubic model and $R^{2}$ = 0.11 with a linear model), is as much as three times higher in $R^{2}$ values than the other correlations.
@@ 3 Discussion
Across all cities chosen, the statistical analysis revealed that differences in global urban visual perception are explained by demographics and personalities (Figure 2). While existing work found safety perceptions, in specific locations, are explained by gender [34, 35, 15, 14, 16], our findings revealed a potential aggregated difference when considering all five locations together. Similar to work done in [11], we did not find statistically significant differences between age groups for safety perceptions, although the age group representation in said work was less granular: below or above 50 years old. Another work also did not find significant differences in perception scores by age or gender [9], although similarly to [11], age groups were analyzed as below or above a given age, i.e., 28 years old in said study. Unlike work in [34], we did not find significant differences in the boring indicator for gender, or any, demographic subgroups. Nevertheless, our findings show that in other indicators, both traditional and new, there are significant differences for all five demographic characteristics in different countries. More importantly, these significant differences are found in cities from continents often overlooked by the research community, e.g., Africa [36]. Furthermore, although nested demographic groups presented limitations in sample size, we found significant differences with reasonable sample sizes ($n\geq 20$) in groups that included AHI (Figure 9).
@@
Our findings on personality traits follow this same trend (Figure 3). With a similar amount of differences compared to those observed across single demographic subgroups, Abuja (Nigeria), Singapore, and San Francisco and Santa Clara (USA) exhibit differences among residents with distinct personality traits, e.g., personality scores in either Q1 or Q3 of their cities. Personality traits have been linked with subjective well-being, for example, conscientiousness was found to be a good predictor of life satisfaction [37]. Personality subgroups based on this, and the extraversion trait, led to the most significant differences including when all participants’ ratings were combined (middle plot in Figure 3). These results open the opportunity to quantify how urban streetscapes are perceived by residents of different levels of well-being. Despite the remaining three personality traits showing only one significant association, our overall findings reinforce ample literature correlating personality and landscape preference and the relationship of personality and aesthetic judgment [38]. These links are crucial when designing tailored urban solutions as most of the significant differences over demographic and personality profiles were found within countries, 15 differences compared to four differences across all locations.
@@
Our quantitative study expands the understanding of machine learning model magnitude bias on urban visual perception prediction, corroborates the perception differences from different populations and how different cities are perceived, and quantifies the influence of someone’s location when rating their cities (Figure 4). Supporting previous ideas of cities’ identity and their characteristics being captured by machine learning models [39, 40], we found that the predictions of ViT-PP2, a model trained on PP2 [6], differ consistently with our diverse dataset. While we noted differences between how participants from each city rated all SVI and between how all participants rated each city (top and middle plot in Figure 4), the predicted perception scores paint optimistic and pessimistic scenarios for the different cities. In both scenarios, albeit with some exceptions, the predicted perceptions of a one-size-fits-all and off-the-shelf model are statistically higher than the ground truth scores under positive indicators and statistically lower under negative indicators. Moreover, participants seem to rely on visual familiarity; influenced by cultural differences, location-specific norms, and preconceived conceptions of the geography; when rating their own city and others. Despite not having information on the location of the SVI, i.e., the pairwise comparison did not show the city the SVI is from, participants appeared to either be familiar with their city or could identify characteristics similar to their city and use that as a reference when rating other cities they might not be familiar with (bottom plot in Figure 4).
@@
Our work asserts the importance of multi-city and multi-population analysis. Previous multi-country work either analyzed only a single country of residency of a dataset [9], did not consider participants’ location [6], or focused on a single feature, e.g., biophilia [8]. Equally importantly, our work evidences that targeted intervention should consider the perception of locals in their own city [42], as it will differ greatly if the perceptual evaluation is done using external SVI or by a machine learning model prediction [16]. Nevertheless, we do not claim generalization or universality of these patterns, and we are limited to evaluating a single city in each country as the scope of this work. Our contribution lies in thoroughly, and for the first time, assessing and quantifying these perceptual differences across multiple locations and populations, highlighting the importance of human-centric and context-aware methodologies for future larger urban visual perceptions studies. While we handled inter-country biases by analyzing demeaned perception scores in their respective locations (bottom plot in Figure 4), in-depth analysis on cross-cultural, location-based biases remains a promising research direction. At the same time, intra-country differences should be further analyzed, especially in larger and more diverse countries.
@@
Despite the demographic diversity and geographic location of participants, most of our proposed indicators (live nearby, cycle, and walk) capture, predominantly linearly, the trends of the existing six indicators. These new dimensions are able to expand perception studies with indicators reflecting a user’s willingness to be associated with an urban scene (e.g., live, walk, or cycle in the surroundings). The green indicator introduces the opportunity to expand greenery assessment from traditional objective measurements such as pixel-ratio metrics, e.g., Green View Index (GVI), to subjective context-aware human perception data. While the relationships found with other perceptual indicators were more complex and nonlinear, we found around half of the statistically significant differences in perception scores between demographic and personality profiles (6 out of 13) under these four new indicators. These highlight the potential of these indicators as input for targeted interventions. As different demographic subgroups in diverse location perceive them differently; e.g., different live nearby and cycle preferences between gender groups in San Francisco and Santa Clara (USA), and Abuja (Nigeria), respectively (Figure 2); different green preferences between high and low agreeableness groups in Singapore; and different walk preferences in genders of a given age group in San Francisco and Santa Clara (USA) (Figure 9). These indicators offer a clearer human subjective perspective that, when complemented with objective measurements, leads to improved urban solutions and assessments [42, 11, 33]. We propose that researchers consider including these indicators where possible and when relevant.
@@
This work contributes to the few studies in multi-city perception, emphasizes the importance of human-centric data collection, and releases a global and diverse perception dataset. Methodologically, we introduced a global and, participants-wise, demographically balanced urban visual perception called SPECS and we thoroughly analyzed perception scores by demographic and personality subgroups. Moreover, we quantified and statistically verified, for the first time, the magnitude bias of off-the-shelf machine learning models in perceptual predictions despite being trained on a global dataset like PP2, we isolated and analyzed the effect of how people are influenced by preconceived conceptions of urban contexts, and introduced and validated new indicators relevant to the six commonly used. Our results draws us closer to understanding in what contexts, e.g., locations and perceptual indicators, urban visual perception is most affected by demographics and personality. While we are not able to pinpoint a specific number or percentage of how much people’s profiles drive perception responses, we found that as such it can lead to significant differences depending on the profiles and locations. These insights on profiles and location subgroups can be extended to other urban studies, such as building exterior evaluation for urban development and strategies [43]. Street-level image-based indicators closely correlate with a city’s social structure and economic development [44]. Thus, as SVI captures the human-level perspective, more closely aligned with people’s experiences in cities, studying and understanding urban streetscapes allows researchers and stakeholders to take informed actions [45]. While PP2 remains an impressive effort valuable for urban science, our analysis reveals its limitations in localized contexts, suggesting that generalized or one-size-fits-all models may produce distorted results. We recommend that researchers fine-tune their perception prediction models with local data, even though this may require additional data collection efforts. This approach is likely to produce more accurate and better performing models [11, 16].
@@
Secondly, we acknowledge the limited number of cities and countries chosen as well as the respective number of images. Our approach of one city per country per continent limits the generalizability of our findings to the broader urban landscape worldwide. To mitigate these limitations, we employed stratified sampling of both urban environments (image selection) and participant demographics, ensuring proportional representation of local socioeconomic contexts within each selected city. Additionally, to ensure statistical power across all ten different indicators for demographic and personality subgroups’ perception ratings analyses, we constrained the number of images per city to 80, to collect enough participant ratings per image and per subgroup to support robust comparisons across multiple demographic and personality dimensions. Rather than claiming global or country-level representativeness and insights, our study focuses on determining whether perceptual indicators vary systematically across demographic groups in diverse urban environments. The reduced number of images per city, taken primarily from city centers, might introduce location biases that could affect the interpretation of our findings. These location biases may include overrepresentation of tourist areas, commercial districts, or architecturally distinctive features that may not reflect typical residential or suburban environments within each city, e.g., very unique urban forms that people associate with specific cities, such as Amsterdam’s canals. While our stratified sampling approach helps ensure demographic representativeness within each city, it cannot fully address the geographical limitations of our urban sampling strategy. Nevertheless, our contributions rely on thorough comparisons of perceptions across different demographic subgroups (top and middle plots in Figure 2), with analyses that control for location when possible (bottom plot in Figure 4). This methodological approach allows us to isolate demographic effects on urban perception across multiple diverse urban contexts, i.e., five cities in five different continents compared to perception studies routinely using one or two cities [5, 31, 32, 33].
@@
Future research should prioritize analyzing interactions between multiple demographic factors, e.g., older adults with different walk abilities [46], and their combined effects on various perceptual indicators. As our study nested participant subgroups by demographics, the resulting smaller participant pools yielded fewer ratings per perceptual indicator (e.g., safety perception scores based solely on older women’s ratings). To address this limitation, subsequent studies could either increase participant numbers or ratings per image. The latter could be achieved by reducing the number of images per city, though potentially compromising urban landscape diversity, or by focusing on fewer perceptual indicators, such as exclusively studying safety perception. Another alternative is to leverage scalable perception rating methods. Human-machine frameworks [47] allows for rapid and cost-effective perception assessments. Building on our findings on the demographic and location-specific perceptual differences, these methods could be more appropriately deployed to quickly evaluate entire cities.